what is the solution to the question pls help 
(sin a+sin 2a)/1+cos a+cos 2a=
 

Grade:11

1 Answers

midhun
7 years ago
 we know
Cos 2a=2 cosa-1
2cos2a= cos2a+1
 
 
So
(Sin a+sin2a)/(1+ cos a+ cos 2a)=(sina +2sina.cosa)/(cos a+(1+ cos 2a)
= sin a(1+2 cos a)/(cos a+2 cos2 a)
Let 1+ 2cos=g
For my simplicity.= sin a . G/cos a .  G
= tan a